Joint Cost Allocation—Weighted Average Method

Gary’s Grooves Co. produces two types of carving knives, one with a handle made of a polymer that looks like walnut wood and another with a handle made with a polymer that looks like red oak. The knives are made through a joint production molding process that produces 330 knife blades for red oak handle knives and 220 knife blades for walnut handle knives at the split-off point. The polymer for the red oak handle knife blades requires twice as much cooling time as the polymer for the walnut handle knife blades, although all knives are removed from the joint molding process at the same time (i.e., once the cooling for the red oak handle knives is complete). The joint production process costs a total of $6,500.

Assuming the company allocates joint costs using the weighted average method based on the required cooling time of the two joint products, determine the amount of joint production costs allocated to each type of knife using the weighted average method.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Thanks!
Gary’s Grooves Red Oak Walnut Total Note
Units Produced      330.00       220.00 A
Weights assigned           2.00           1.00 B
Weighted production units      660.00       220.00       880.00 C=A*B
Joint costs 6,500.00 D
Cost per Weighted units           7.39 E=D/C
Cost per units         14.77           7.39 F=E*B
Cost allocated 4,875.00 1,625.00 6,500.00 G=F*A